Home Clean room MedAustron HEPHY testbeams old
testbeam-DESY-Oct2017 testbeam-DESY-April2017 testbeam-SPS2015 testbeam-SPS2014 testbeam-DESY14 testbeam-SPS12 testbeam-SPS11 testbeam-SPS10 testbeam-SPS09 testbeam-SPS08 LP-TPC
  Alibava/Mapsa testbeam at DESY in April 2017 (Axel, Grossi)  Not logged in ELOG logo
Message ID: 67     Entry time: Fri Apr 14 14:21:35 2017
Author: Axel Koenig 
Project: common 
Subject: Run summary IFX_CB  
Run Number:  
Events:  
XYPosition:  
Cooling: No 
Data: Unknown 
Alibava run Telescope run File existence check Run type Triggers Remarks
001 496 yes position check 200000 none
002 498 yes beam focused on one strip segment only 100000 none
003 500 yes beam focused on one strip segment only 100000 none
005 508 yes Bias scan. Voltage = 10 V 285000 none
006 510 yes Bias scan. Voltage = 20 V 185000 none
007 513 yes Bias scan. Voltage = 30 V 170000 none
008 515 yes Bias scan. Voltage = 40 V 175000 none
009 517 yes Bias scan. Voltage = 50 V 175000 none
010 521 yes Bias scan. Voltage = 60 V 175000 none
011 522 yes Bias scan. Voltage = 65 V 175000 none
012 523 yes Bias scan. Voltage = 70 V 175000 none
013 524 yes Bias scan. Voltage = 75 V 175000 none
014 525 yes Bias scan. Voltage = 80 V 175000 none
015 526 yes Bias scan. Voltage = 90 V 175000 none
016 527-531 yes Bias scan. Voltage = 100 V 750000 Telescope files 527 to 531, run crashed!
017 532 yes Bias scan. Voltage = 110 V 175000 none
018 533 yes Bias scan. Voltage = 120 V 175000 none
020 535 yes overnight run at 100 V 190000 run crashed data may be corrupted
021 540 yes CRAP! ? CRAP!
022 545 yes beam focused on one strip segment only 500000 none
023 546 yes beam focused on one strip segment only 385000 run crashed data may be corrupted
024 547 yes Bias scan. Voltage = 140 V 175000 none
025 548 yes Bias scan. Voltage = 150 V 175000 none
026 549 yes Rotation -10 deg 350000 none
027 550 yes Rotation +10 deg 350000 none
028 551 yes overnight run at 100 V 3000000 run crashed data may be corrupted
029 557 yes Bias scan. Voltage = 130 V 175000 none
030 558 yes long run 1000000 telescope chiller turned off (keept at this state unless otherwisde stated

Central beam position: x = 109 mm, y = 4.5 mm, rot = 240 deg

Current was stable for all runs i.e. 140 +- 10 nA at 100 V

Trigger efficiency ~ 60% for all runs (except bias scan).

Temperature in area 20+-1 deg.

ELOG V3.1.4-966e3dd